40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 8 20:11:54 2012, Jenne, Update, Alignment, AS steering moved out of POY's way, 2 green beams onto PSL table AS_8Nov2012.pngREFL_8Nov2012.png
    Reply  Thu Nov 8 20:54:08 2012, Manasa, Ayaka, Update, Alignment, Reconfirming on IPPOS, IPANG and oplevs centering 
    Reply  Fri Nov 9 08:11:57 2012, Steve, Update, Alignment, nothing moved - all good opsingGOOD.pngop8h.png8amF.pngkeepfailing.png
       Reply  Fri Nov 9 11:38:38 2012, jamie, Update, General, we're closing up 
          Reply  Fri Nov 9 17:15:05 2012, Manasa, Steve, Ayaka, Update, General, We're closed! Pumping down monday morning 
Message ID: 7690     Entry time: Thu Nov 8 20:54:08 2012     In reply to: 7689
Author: Manasa, Ayaka 
Type: Update 
Category: Alignment 
Subject: Reconfirming on IPPOS, IPANG and oplevs centering 

Quote:

" We found that IPANG was not on its photodiode, but determined that it was centered on all of the in-vac mirrors, and that it was just a little bit of steering on the ETMY end out-of-vac table that needed to be done."

Manasa took photos of all test mass chambers and the BS chamber, so we can keep up-to-date CAD drawings. 

Oplevs and IPPOS/IPANG are being centered as I type.  Manasa and Ayaka are moving the lens in front of IPANG such that we have a slightly larger beam on the QPD.

 

The lens in front of IPANG on the out-of-vac table was moved to get a larger beam giving reasonable signals at the QPD.

IPPOS did not need much adjustment and was happy at the center of the QPD.

All oplevs but the ETMY were close to the center. I had to move the first steering mirror about half an inch on the out-of-vac table to catch the returning oplev beam from ETMY and direct it to the oplev PD.

* We have taken reasonable amount of in-vac pictures of ETM, ITM and BS chambers to update the CAD drawing.

 

ELOG V3.1.3-